# Revisiting Multi-Agent World Modeling from a Diffusion-Inspired Perspective

### v1

@@ Appendix A Proof of Theorem 2
=\>= $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k)}|s_{t+1}^{(k-1)})}\right]$
=\>= $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\frac{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})\hat{q}(s_{t+1}^{(k)}|s_{t+1}^{(0)})}{q(s_{t+1}^{(k-1)}|s_{t+1}^{(0)})}}\right]$
=\>= $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\frac{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}{\hat{q}(s_{t+1}^{(n)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})}\right]$
=\>= $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(n)}|s_{t+1}^{(0)})}+\sum_{k=2}^{n}\log\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})}\right]$
@@ Appendix E Additional Results
E.1 Final Returns of All Methods on MAMuJoCo and Bi-DexHands
@@
Walker2d-2x3 $\mathbf{3329}_{\pm 1056}$ $\underline{2822}_{\pm 997}$ $124_{\pm 19}$ $1135_{\pm 210}$ $1007_{\pm 282}$ $752_{\pm 216}$
@@
BottleCap 300K $\mathbf{259.9}_{\pm 4.1}$ $203.8_{\pm 5.2}$ $\underline{210.9}_{\pm 6.1}$ $100.7_{\pm 3.8}$ $104.0_{\pm 2.3}$
DoorOpenInward $\mathbf{290.4}_{\pm 29.0}$ $225.0_{\pm 79.4}$ $\underline{246.3}_{\pm 7.0}$ $30.7_{\pm 2.5}$ $65.8_{\pm 6.9}$
DoorOpenOutward $\mathbf{367.1}_{\pm 19.4}$ $177.4_{\pm 43.1}$ $\underline{221.9}_{\pm 7.3}$ $58.8_{\pm 4.6}$ $96.4_{\pm 8.5}$
BottleCap $\mathbf{24.4}_{\pm 11.4}$ $\underline{4.3}_{\pm 0.4}$ $0.0_{\pm 0.0}$ $0.0_{\pm 0.0}$ $0.0_{\pm 0.0}$
E.2 Experiment Details of Imagination Evaluation across Different Conditioning Orders
@@ G.4 DIMA hyperparameters
Number of denoising steps {$&2\cdot|\mathcal{N}|\text{ if }|\mathcal{N}|\leq 2\\ &|\mathcal{N}|\text{ if }|\mathcal{N}|>2\end{aligned}\right.$
Procedure training_loop():
       for epochs do
             collect_experience(steps_collect)
             for steps_state_decoder do
                   update_state_decoder()
            for steps_diffusion_model do
                   update_diffusion_model()
            for steps_reward_end_model do
                   update_reward_end_model()
            for steps_actor_critic do
                   update_actor_critic()
Procedure collect_experience($n$):
@@
       for $t=0$ to $n-1$ do
             Sample $a_{t}^{i}\sim\pi_{\psi}^{i}(a_{t}^{i}|o_{t}^{i}),\,\forall\text{ agent }i$
             if $\gamma_{t}=1$ then
[…]
Procedure update_state_decoder():
       Sample state-observation pair $(s_{t},o_{t}^{1:n})\sim\mathcal{D}$
       Compute $\mathcal{L}_{\text{FSQ}}$ in Eq. (22)
       Update State Decoder $g_{\varphi}$
Procedure update_diffusion_model():
       Sample $s_{t+1}^{\tau}\sim\mathcal{N}(x_{t+1}^{0},\sigma^{2}\mathbf{I})$
       Sample a chosen agent id $k\sim\text{Uniform}\{1,2,\dots,n\}$
       Update Diffusion Model $D_{\theta}$
Procedure update_reward_end_model():
       for $i=t$ to $t+H-1$ do
       Update Reward and Termination Model $f_{\phi}$
Procedure update_actor_critic():
       Set the joint action condition order $\rho=(i_{1},i_{2},\dots,i_{n})$
       Let $\hat{o}_{t}^{1:n}=o_{t}^{1:n}$
       for $i=t$ to $t+H-1$ do
Algorithm 1 DIMA paired with learning in imaginations

### v2

@@ Appendix A Proof of Theorem 2
$=\>$ $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k)}|s_{t+1}^{(k-1)})}\right]$
$=\>$ $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\frac{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})\hat{q}(s_{t+1}^{(k)}|s_{t+1}^{(0)})}{q(s_{t+1}^{(k-1)}|s_{t+1}^{(0)})}}\right]$
$=\>$ $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}+\log\frac{\hat{q}(s_{t+1}^{(1)}|s_{t+1}^{(0)})}{\hat{q}(s_{t+1}^{(n)}|s_{t+1}^{(0)})}+\log\prod_{k=2}^{n}\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})}\right]$
$=\>$ $\mathbb{E}_{\hat{q}(s_{t+1}^{(1):(n)}|s_{t+1})}\left[\log\frac{p(s_{t+1}^{(n)})p(s_{t+1}^{(0)}|s_{t+1}^{(1)},s_{t},a_{t}^{1})}{\hat{q}(s_{t+1}^{(n)}|s_{t+1}^{(0)})}+\sum_{k=2}^{n}\log\frac{p(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t},a_{t}^{k})}{\hat{q}(s_{t+1}^{(k-1)}|s_{t+1}^{(k)},s_{t+1}^{(0)})}\right]$
@@ Appendix E Additional Results
E.1 Additional Experiments: Detailed Returns of All Methods on MAMuJoCo and Bi-DexHands
@@
Walker2d-2x3 $\mathbf{3329}_{\pm 1056}$ $\underline{2822}_{\pm 997}$ $24_{\pm 19} 11$$135_{\pm 210}$ $1007_{\pm 282}$ $752_{\pm 216}$
@@
BottleCap 300K $\mathbf{259.9}_{\pm 4.1}$ - $203.8_{\pm 5.2}$ $\underline{210.9}_{\pm 6.1}$ $100.7_{\pm 3.8}$ $104.0_{\pm 2.3}$
DoorOpenInward $\mathbf{290.4}_{\pm 29.0}$ - $225.0_{\pm 79.4}$ $\underline{246.3}_{\pm 7.0}$ $30.7_{\pm 2.5}$ $65.8_{\pm 6.9}$
DoorOpenOutward $\mathbf{367.1}_{\pm 19.4}$ - $177.4_{\pm 43.1}$ $\underline{221.9}_{\pm 7.3}$ $58.8_{\pm 4.6}$ $96.4_{\pm 8.5}$
BottleCap $\mathbf{24.4}_{\pm 11.4}$ - $\underline{4.3}_{\pm 0.4}$ $_{\pm 0.0} 0.0±0.0$$_{\pm 0.0}$ $0.0_{\pm 0.0}$
Table 6: Policy learning performance (final return) of DIMA with different imagination horizons ($H=15$ vs. $H=25$).
Scenarios DIMA w/ $H=15$ DIMA w/ $H=25$
Ant [4x2] (4-agent) $\mathbf{4766}_{\pm 450}$ $4328_{\pm 1058}$
HalfCheetah [6x1] (6-agent) $5643_{\pm 163}$ $\mathbf{6310}_{\pm 335}$
E.2 Additional Experiments: Policy Learning with Longer Imagination Horizons
We investigated the downstream impact on policy learning when utilizing the extended imagination horizon ($H=25$) during the whole training process. The final policy returns are presented in Table 6. The results are environment-dependent and provide critical insights. On HalfCheetah [6x1], training with a longer horizon ($H=25$) yields a substantial performance improvement. This strongly suggests that when DIMA’s long-range predictions are stable, the policy optimizer can successfully leverage these extended rollouts to discover more complex and far-sighted strategies, leading to superior returns. Conversely, on Ant [4x2], we observe a performance degradation accompanied by significantly higher variance. This indicates that for this specific task, while comparatively lower than baselines (as shown in Table 1), the accumulated prediction errors in $H=25$ rollouts are still sufficient to introduce noise that misleads the policy learning. This phenomenon underscores both the practical potential of DIMA in scalable model-based MARL (demonstrated on HalfCheetah) and highlights a key challenge: ensuring that the predictive accuracy is robust enough to provide a stable gradient signal for policy learning across all task types.
E.3 Additional Experiments: Comparison on Training Compute
We measure the training time and GPU memory usage of all evaluated model-based MARL methods, including our proposed DIMA. All experiments were conducted using a single NVIDIA RTX 4090 GPU to ensure a fair comparison across methods and tasks. As shown in the table below, DIMA is substantially more efficient than MARIE in both training time and GPU memory usage, and is comparable to or slightly more efficient than MAMBA.
Table 7: Comparison on consumed computational resources over 2 test scenarios.
Methods Training Time Usage of GPU Mem
Ant-2x4
DIMA  1d 19h 3.10 GB
MARIE  3d 17h 14.33 GB
MAMBA  1d 1h 2.74 GB
Ant-4x2
DIMA  1d 19h 3.15 GB
MARIE  3d 17h 3.10 GB
MAMBA  1d 1h 4.37 GB
E.4 Ablation: Sequential vs. Joint Modeling Prediction Accuracy
Table 8: Ablation study comparing the cumulative L1 observation errors of sequential vs. joint modeling. Models were trained on 500k transitions and evaluated on a 500k held-out set. Sequential modeling achieves statistically indistinguishable prediction accuracy, validating its design.
Task Method Obs L1 Error @ $H=15$ Obs L1 Error @ $H=20$
DoorOpenOutward Sequential (Ours) $\mathbf{5.333}_{\pm 0.273}$ $\mathbf{7.081}_{\pm 0.325}$
Joint ${5.345}_{\pm 0.267}$ ${7.092}_{\pm 0.324}$
DoorOpenInward Sequential (Ours) $\mathbf{5.563}_{\pm 0.326}$ $\mathbf{7.447}_{\pm 0.393}$
Joint ${5.565}_{\pm 0.322}$ ${7.453}_{\pm 0.386}$
Pen Sequential (Ours) $\mathbf{6.667}_{\pm 1.764}$ $\mathbf{8.936}_{\pm 2.328}$
Joint ${6.676}_{\pm 1.762}$ ${8.947}_{\pm 2.322}$
E.5 Experiment Details: Imagination Evaluation across Different Conditioning Orders
@@ Appendix F Overview of DIMA with Learning in Imaginations
Procedure training_loop():
 for epochs do
    collect_experience(steps_collect)
    for steps_state_decoder do
       update_state_decoder()
    for steps_diffusion_model do
       update_diffusion_model()
    for steps_reward_end_model do
       update_reward_end_model()
    for steps_actor_critic do
       update_actor_critic()
Procedure collect_experience($n$):
 $s_{0},o_{0}^{1:n}\leftarrow\texttt{env.reset()}$
 for $t=0$ to n1n-1 do
      Sample $a_{t}^{i}\sim\pi_{\psi}^{i}(a_{t}^{i}|o_{t}^{i}),\,\forall\text{ agent }i$
    if $\gamma_{t}=1$ then
       $s_{t+1},o_{t+1}^{1:n}\leftarrow\texttt{env.reset()}$
Procedure update_state_decoder():
   Sample state-observation pair $(s_{t},o_{t}^{1:n})\sim\mathcal{D}$
   Compute $\mathcal{L}_{\text{FSQ}}$ in Eq. (22)
   Update State Decoder $g_{\varphi}$
Procedure update_diffusion_model():
   Sample $s_{t+1}^{\tau}\sim\mathcal{N}(x_{t+1}^{0},\sigma^{2}\mathbf{I})$
   Sample a chosen agent id $k\sim\text{Uniform}\{1,2,\dots,n\}$
   Update Diffusion Model $D_{\theta}$
Procedure update_reward_end_model():
 for $i=t$ to t+H1t+H-1 do
   Update Reward and Termination Model $f_{\phi}$
Procedure update_actor_critic():
   Set the joint action condition order $\rho=(i_{1},i_{2},\dots,i_{n})$
   Let $\hat{o}_{t}^{1:n}=o_{t}^{1:n}$
 for $i=t$ to t+H1t+H-1 do
Algorithm 1 DIMA paired with learning in imaginations
@@ G.4 DIMA hyperparameters
Appendix H Broader Impact
Our work introduces DIMA, a diffusion-inspired multi-agent world model that significantly improves sample efficiency in cooperative multi-agent control environments. By enabling more faithful imagined rollouts, DIMA can accelerate the development of complex autonomous systems—such as multiple real robots coordination, traffic management, and energy-efficient buildings—thereby reducing real-world trial costs. However, these capabilities also carry potential risks: misuse of high-fidelity neural simulators for adversarial planning could exacerbate privacy and security concerns.